OpenStudy (anonymous):

solve the following system of equations x+3y = 5 (1) x = 4 - 3y (2) what is the solution of the system? I got an answer of no solution is that correct?

OpenStudy (amistre64):

x+3y = 5 x = 4 - 3y [-1] x+3y = 5 -x = -4 + 3y --------------- 3y = 1+3y .... i wonder lol

OpenStudy (amistre64):

they are parallel lines that never meet

OpenStudy (saifoo.khan):

x = 4.5 y = 1/6

OpenStudy (asadkarim7):

ES calculator saifoo ?

OpenStudy (saifoo.khan):

haha, Yeah! :D

OpenStudy (anonymous):

so is there a solution or not?

OpenStudy (anonymous):

hi amistre64 so is there a solution or not?

OpenStudy (amistre64):

a solution would imply that the lines cross at some point. no solution implies that they lines will never cross. and infinite solutions implies that the lines are the same line.

OpenStudy (amistre64):

since these lines never meet .....

OpenStudy (anonymous):

so this means there is no solution without the lines meeting

OpenStudy (amistre64):

that is correct :)

OpenStudy (anonymous):

thanks amistre64 can i become your fan

OpenStudy (amistre64):

if you want :)

OpenStudy (anonymous):

yes i like your style